A billiard ball moving at 5.0m/s strikes a stationary ball of the same mass.  After the collision, the ball moves at 4.33m/s at an angle of 30.0° with respect to the original line of motion.  Assuming an elastic collision, find the magnitude and direction of the struck ball’s velocity.



1
Expert's answer
2021-05-26T10:32:16-0400
"v_{2x}'=\frac{5}{4.33\cos 30}=1.25\frac{m}{s}\\v_{2y}'=-4.33\sin{30}=-2.16\frac{m}{s}"

"v_2'=2.5\frac{m}{s}\\\theta=300\degree"